### v4

# Universal Approximation Theory: Foundations for Parallelism in Neural Networks

Abstract
Neural networks are increasingly evolving towards training large models with big data, a method that has demonstrated superior performance across many tasks. However, this approach introduces an urgent problem: current deep learning models are serial computing, meaning that as the number of network layers increases, so do the training and inference times. This is unacceptable if deep learning models go deeper. Therefore, this paper proposes a deep learning parallelization strategy based on the Universal Approximation Theorem (UAT). From this foundation, we designed a parallel network called Para-Former to test our theory. Unlike traditional serial models, the inference time of Para-Former does not increase with the number of layers, significantly accelerating the inference speed of multi-layer networks. Experimental results validate the effectiveness of this network.
1 Introduction
Currently, deep learning is advancing towards larger models and datasets, driven by the success of ChatGPT (Radford and Narasimhan 2018; Brown et al. 2020; Achiam et al. 2023), Llama (Touvron et al. 2023), and ViT (Dosovitskiy et al. 2020), among others. These large models demonstrate significantly enhanced performance compared to their predecessors. According to UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024b), residual-based convolutional neural networks (CNNs) and Transformer-based large language models are specific implementations of the UAT (Cybenko 2007; Hornik, Stinchcombe, and White 1989), capable of dynamically approximating functions based on inputs. From the perspective of UAT, larger models inherently possess stronger dynamic fitting capabilities (Wang and Li 2024b, a). Therefore, the trend towards larger models is undoubtedly the future direction of deep learning, and large datasets are essential for training these models to ensure their approximations accurately reflect real-world complexities.
However, this trend towards larger models and datasets inevitably leads to increased demands for computational resources and extended training times. Despite advancements in hardware technology, such as GPUs and TPUs, the rise in time costs remains a significant obstacle. The speed of electrical signal transmission is limited, whereas the theoretical scale of networks has no such upper limit. As network layers increase to tens or even hundreds of thousands, inference times can extend significantly, potentially requiring several minutes or hours for a single batch, making training large models on large datasets exceedingly difficult.
Current solutions include model optimization techniques such as quantization (Jacob et al. 2017; Lu, Zhong, and Yang 2022; Guo, Wang, and Cui 2021; Lin et al. 2023), pruning (Ma, Fang, and Wang 2023; Xia et al. 2023), and knowledge distillation (Gou et al. 2020; Gu et al. 2021). Additionally, specialized accelerators (Jouppi et al. 2017; Elbtity, Chandarana, and Zand 2024) and edge computing (Shi et al. 2016; Gale et al. 2020), as well as parallel and distributed computing (Shallue et al. 2018; Ben-Nun and Hoefler 2018; Sergeev and Balso 2018), offer potential solutions. However, these methods do not fundamentally address the inference delay caused by the increasing number of serial network layers.
The root cause lies in the serial computation mechanism of existing deep learning models, where the output of one layer serves as the input for the next. This design originated from early computer vision networks, which used convolutional layers with small receptive fields to accommodate the local correlations and spatiotemporal invariance of image data. Achieving a larger receptive field required deeper network layers, as exemplified by architectures like AlexNet (Krizhevsky, Sutskever, and Hinton 2012) and VGG (Simonyan and Zisserman 2014). The introduction of ResNe t (He et al. 2015) marked a paradigm shift, as the residual structure could dynamically fit functions based on input (Wang and Li 2024b, a), influencing subsequent network designs and establishing the tradition of multi-layer stacking. While this design was manageable with fewer layers, the increasing depth and number of parameters in modern networks have created a critical issue: slow computation efficiency due to the need for sequential layer-by-layer calculations during inference. This computation method inevitably leads to delays, which become more severe as network depth increases, significantly constraining the progress of computer vision and natural language processing technologies.
Therefore, it is imperative to explore parallelization techniques for deep learning networks, enabling parallel computation across multiple layers. This approach is key to accelerating the inference process, breaking the time bottleneck, and advancing computer vision and natural language processing technologies towards greater efficiency and speed. To develop such parallel computation, we must start from the fundamental theory of deep learning: the UAT. The network must be designed in accordance with UAT principles and must satisfy the requirements outlined in the papers UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024a), which emphasize the ability to dynamically fit functions based on data. With this in mind and leveraging existing network structures, we designed a parallel network, Para-Former, to validate the feasibility of our theory. Our contributions include:
  • Proposing the theory of parallel deep learning networks and experimentally validating its feasibility.
  • Analyzing the characteristics, advantages, and disadvantages of both serial and parallel networks.
  • Investigating the issues present in deep learning from a data perspective.
The structure of this paper is as follows: In Section 2, we first present the general form of the UAT. Based on the mathematical form of UAT, we define the parallel network Para-Former in Section 3. In Section 4, we validate the effectiveness of Para-Former in various ways: we assess its feasibility in Section 4.2, explore the impact of network depth in Section 4.3, and analyze the influence of datasets on network performances in Section 4.4.
2 UAT
Deep learning networks have demonstrated their powerful capabilities in numerous CV and NLP tasks. In the papers UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024a), they have been proven that both Transformer-based networks and residual-based CNNs are specific implementations of the UAT. Therefore, we consider UAT to be the fundamental theory for deep learning. To design parallelized networks, we must start from the principles of UAT. Before formally designing parallel networks, we will first review the basic form of the UAT theory and then propose a network parallelization solution based on this mathematical theory. The UAT was initially proposed by Cybenko (Cybenko 2007), and its basic form is as follows:
$G(\mathbf{x})=\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (1)
is dense in $C\left(\mathbf{I}_{n}\right)$. Here, $\mathbf{W}_{j}\in\mathbb{R}^{n}$ and $\alpha_{j},\theta\in\mathbb{R}$ are parameters $\sigma$ is a sigmoid function. For any $f\in C\left(\mathbf{I}_{n}\right)$ and $\varepsilon>0$, there exists a function $G(\mathbf{x})$:
$|G(\mathbf{x})-f(\mathbf{x})|<\varepsilon\quad\text{ for all }\quad\mathbf{x}\in\mathbf{I}_{n}.$ (2)
This indicates that when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik, Stinchcombe, and White  (Hornik, Stinchcombe, and White 1989) further proved that this function can approximate any Borel measurable function. Observing Eq. (1), where the function $G(\mathbf{x})$ produces scalar outputs in the real number field $\mathbb{R}$, the scenario naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation in each dimension. To accommodate such multi-dimensional outputs, we simply need to extend Eq. (1): the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix. We call the parameters $\mathbf{W}_{j}$ and $\alpha_{j}$ in UAT as weight and $\theta_{j}$ as bias.
Figure 1: The general description of the serial network and parallel network.
3 UAT for Parallel Network
3.1 Requirements for Basic Block in Parallel Network
From Eq. (1), we can observe a significant characteristic: if we define different indices $j$ as the indices of layers and the corresponding layer’s mathematical form is $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$, it is evident that the parameters between different layers do not interact with each other. Therefore, based on Eq. (1), we can design networks where parameters across multiple layers are independent, achieving parallel computation, like Figure 1. Figure 1.a represents a typical sequential network, which clearly requires computation to proceed in order from front to back. In contrast, Figure 1.b depicts a parallel network, where there is no parameter interaction between different blocks, allowing them to be considered independent and thus suitable for parallel computation. The blocks in this network should satisfy the formula of $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$. Clearly, under this condition, the sum of the multiple blocks’ outputs adheres to the UAT theory.
If we design the parallel network solely based on Eq. (1), the network will face a limitation once training is complete: the function that can be approximated by the network is fixed, as the parameters in the corresponding UAT are fixed. So to ensure the network can dynamically approximate functions based on input, some or all of the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ must be influenced by the input (referencing to UAT2LLMs and UAT2CVs), which can be simply written as $\alpha_{j}(\mathbf{x})$, $W_{j}(\mathbf{x})$ and $\theta_{j}(\mathbf{x})$. This allows the network to dynamically adapt and fit the function based on the input while enabling parallel computation. So we could design the basic block in Figure 1, like $\text{Block}_{1}\cdots\text{Block}_{n}$ are the implementation of $\alpha_{j}(\mathbf{x})\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}(\mathbf{x})\mathbf{x}+\theta_{j}(\mathbf{x})\right),i=1\cdots n$. Some or all the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ will change according to input.
Figure 2: The process of Transformer module.
Figure 3: The structure of the Para-Former-1-n represents Para-Former with 1 depth of n layers.
Figure 4: The structure of the Para-Former-m-n represents Para-Former with m depth of n layers.
Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
Classes 10 100 10 37 102
Train set 47500 47500 4750 3,515 969
Val set 2500 47500 250 185 51
Test set 10000 10000 800 3669 6149
Per class 4750 4750 475 95 9.5
Table 1: This table presents the sizes of the training, validation, and test sets for the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102. It also provides the number of categories (Classes) and the average number of images per category for each dataset (Per class).
3.2 Para-former
Based on the previous request on the basic module, we designed the Para-Former which is a parallel network by using the Transformer module as the basic module. We chose Transformer for this design because it is a well-established and widely-used network architecture, excelling in both CV and NLP. Another reason is that its basic mathematical form aligns with the requirements of Section 3.1.
Before formally introducing the structure of Para-Former, let’s briefly review the Transformer structure. The basic structure of a Transformer is shown in Figure 2, and for convenience, we refer to it as a Transformer module. Typically, it is represented as follows:
$\mathring{\mathbf{H}}=$ $\operatorname{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})$ (3)
$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{i1},\ldots,\hat{\mathbf{H}}_{ih}\right)\mathbf{W}_{O}$
$\hat{\mathbf{H}}_{ij}$ $=\operatorname{Attention}(\mathbf{x}_{ij}\mathbf{W}_{Qij}=\mathbf{Q}_{ij},\mathbf{x}_{ij}\mathbf{W}_{Kij}=\mathbf{K}_{ij},$ (4)
$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{ij})$
$=\operatorname{softmax}\left(\frac{\mathbf{Q}_{ij}\mathbf{K}_{ij}^{T}}{\sqrt{M}}\right)\mathbf{V}_{ij}$
$=\mathbf{H}_{ij}\mathbf{V}_{ij}=\mathbf{H}_{ij}[\mathbf{x}_{ij}\mathbf{W}_{Vij}]$
where the subscript of $ij$ represents the $j$-th head of Multi-Head Attention (MHA) in the $i$-th layer. Please note that throughout the following text, all the subscript $i$ refer to the layer index in the network, like $\mathbf{x}_{i}$. According to Figure 2, assuming the input to the Feed-Forward Network (FFN) in the $i$-th layer of the Transformer, it can be expressed as:
$FFN(\mathbf{x}_{i})=\mathbf{W}_{i,3}\sigma(\mathbf{W}_{i,2}\mathbf{x}_{i}+\mathbf{b}_{i,2})+\mathbf{b}_{i,3}\$ (5)
The above representation mainly describes the engineering process rather than its true mathematical form. According to UAT2LLMs, the MHA and FFN components of Transformers can be expressed in matrix-vector form. We denote the matrix-vector form of variables and operations with a prime symbol (e.g., $\mathbf{x}_{i}$ becomes $\mathbf{x}^{\prime}_{i}$). So MHA and FFN can be written as (For more details, refer to UAT2LLMs):
$MHA^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (6)
$FFN^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (7)
According to Eq. (6) and Eq. (7), the output $\mathbf{x}^{\prime}_{1}$ of a single-layer Transformer can be expressed as Eq. (8) (see Appendix A for more details). The mathematical form of Eq. (8) can be considered equivalent to the two terms $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ in Eq. (1). Therefore, if we parallelize multiple instances of $\mathbf{x}^{\prime}_{1}$, we obtain a multi-layer parallel Transformer network, as illustrated in Figure. 3.
$\mathbf{x}^{\prime}_{1}$ $=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,2})$ (8)
Additionally, we provide the output for a two-layer Transformer, as shown in Eq. (9) (see Appendix A for more details). From Eq. (9), it is evident that the serial configuration of two Transformer layers also adheres to the mathematical form of UAT. Therefore, we can extend the Para-Former structure. Specifically, a block in Figure. 1 can contain multiple serial Transformer modules (it has been proven in UAT2LLMs that the mathematical form of multi-layer Transformers is also consistent with UAT), while multiple blocks operate in parallel, as illustrated in Figure. 4. The number of Transformer modules within a single block is referred to as the depth, while the total number of Transformer modules across all blocks is referred to as the network’s layer count.
$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (9)
$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
Both design strategies possess the capability to dynamically adjust model parameters based on the input. In Figure 3, the primary mechanism is the dynamic adjustment of MHA weight parameters based on the input. In contrast, Figure 4 not only leverages the dynamic adjustment of MHA weight parameters but also utilizes UAT to dynamically approximate some bias parameters. For instance, in Eq. (9), $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$, can be seen as a two-layer UAT. Refer to Appendix B to see which parameters are influenced by the input. So, what are advantages of these strategies? We will talk about these in Section 4.3.
Para-F-1-1 Para-F-1-6 Para-F-1-12 Para-F-1-24
Cifar-10 55.95 64.90 66.88 71.53
Cifar-100 30.67 38.84 41.41 46.96
STL-10 44.41 51.57 55.96 59.53
OxfordIIITPet 19.62 20.95 22.07 22.73
Flower-102 27.33 32.65 37.42 35.84
Table 2: This table shows the predictions of Para-former-1-1, Para-former-1-6, Para-former-1-12, Para-former-1-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102. We simply represent them as Para-F-1-1, Para-F-1-6, Para-F-1-12 and Para-F-1-24.
3.3 Speed-up Ratio
We have proposed a network parallelization scheme, and in this section, we will compare the running speed of this parallelized network with that of traditional serial networks. It is important to note that due to differences in hardware and implementation methods, providing an objective comparison through experiments can be challenging. Therefore, we will focus on a theoretical analysis. Based on the network structure shown in Figure 3 and 4, the inference time of Para-Former is only affected by its network depth.
Assume we use a Para-Former with a depth of $M$ and a layer count of $L$, compared to a traditional Transformer network with a depth of $N$ layers, where $L>>N$. In this case, the running speed of the parallel network would be $N/M$ times faster than that of the serial network. This is because the speed of the parallel network is solely related to the depth of the Para-Former. As the serial network depth increases, the acceleration effect becomes more pronounced.
4 Experiments
Our experimental approach is to first verify the feasibility of the parallel network theory and its adherence to the UAT approximation properties (The larger the value of $N$, the more powerful the model becomes, resulting in better classification performance.) using several commonly used public image classification datasets (Section 4.2). We then explore the impact of network depth on the Para-Former performance (Section 4.3), followed by an investigation into the influence of the dataset (Section 4.4). We conducted experiments with various Para-Former models, where the naming convention uses the first number to indicate the depth and the second number to indicate the number of layers. For instance, Para-former-1-12 denotes a model with 1 depth and 12 layers.
4.1 Dataset
In pursuit of a comprehensive understanding of the efficacy and practicality of parallel models, we meticulously selected several widely-acknowledged and representative image classification datasets for an integrated comparative analysis. These datasets encompass a diverse array of visual content and span various levels of difficulty, thereby furnishing us with a fertile ground to probe the performance boundaries of parallel models. Specifically, our study centers on the following pivotal datasets: CIFAR-10, CIFAR-100 (Krizhevsky 2009), STL-10 (Coates, Ng, and Lee 2011), Flowers-102 (Nilsback and Zisserman 2008) and Oxford-IIIT Pets (Parkhi et al. 2012). In Table 1, we provide more details for these datasets.
Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
Cifar-10 79.39 74.75 77.21 79.92
Cifar-100 55.18 51.04 52.65 56.15
STL-10 62.81 61.60 62.75 64.30
OxfordIIITPet 26.05 23.65 26.02 25.92
Flower-102 37.97  39.17 40.12 40.51
Table 3: This table shows the predictions of Para-former-6-6, Para-former-2-24, Para-former-3-24, Para-former-6-24 on the datasets CIFAR-10, CIFAR-100, STL-10, OxfordIIITPet, and Flower-102.
4.2 The Feasibility of Para-Former
To validate the feasibility and characteristics of Para-Former, we conducted training and validation on several commonly used image classification datasets, with all models trained for 500 epochs. The experimental design involved preprocessing inputs using the common ViT techniques of patch extraction and positional encoding. We then used Transformer as the basic model architecture, setting all network depths to 1 and configuring Para-Former from single-layer to multi-layer setups, as shown in Figure 3 (We evaluated the models’ performance using the accuracy metric, calculated as $\text{accuracy}=100\times\frac{\text{correct}}{\text{total}}$). A single-layer setup corresponds to a ViT with a depth of 1. The results of the models are presented in Table 2.
It is evident from Table 2 that as the number of parallel network layers increases, the model accuracy improves progressively (Except for the result of Para-F-1-24 being worse than Para-F-1-12 on the Flower-102 dataset, this variation is within a reasonable range.). According to UAT theory, increasing the network layers is equivalent to increasing $N$ in Eq. (1), thereby enhancing the approximation capability of UAT. The results in Table 2 are in complete agreement with UAT theory. Therefore, we believe that this parallel network approach is entirely feasible.
However, the overall prediction accuracy in Table 2 is quite low. What could be the reasons for this? We believe there are two main factors affecting the model’s predictive performance:
1. Model Fitting Capability: This is influenced by the model’s size and design, particularly the value of $N$ in the UAT and the degrees of freedom of the parameters. A larger $N$ enhances the UAT’s ability to fit functions. The degrees of freedom refer to the number of parameters in the UAT that are influenced by the input. For instance, the degrees of freedom in Equation (8) is 2, while in Equation (9), it is 6 (see Appendix B). The degrees of freedom represent the model’s ability to dynamically fit corresponding functions based on the input. Simply increasing $N$ can improve the model’s ability to fit specific functions, but images are usually diverse. Thus, the network must have the capability to dynamically fit different functions based on the input. Higher degrees of freedom allow the model to fit multiple different functions according to the input. Additionally, the degrees of freedom vary within the Transformer: they mainly consist of weights and biases. The degrees of freedom for weights stem from the MHA, primarily putting linear transformations based on the input. The degrees of freedom for biases come from residual connections, which can dynamically fit the bias terms using UAT. The deeper the residual layers, the more UAT layers correspond to the bias terms, leading to stronger dynamic fitting capabilities. As a result, a shallow serial network can achieve the fitting ability of a multi-layer parallel network with a depth of 1. We will explore this issue further in Section 4.3.
2. Data Influence: Data is another crucial factor in deep learning training. Since natural images are typically diverse, limited data can lead to overfitting. For example, in a cat-dog classification task, if dogs in the training set mostly appear against backgrounds of blue skies and grass, these backgrounds might be mistakenly considered features of dogs. If cats rarely appear in such backgrounds in the training set but do in the test set, they might be misclassified as dogs. Training on a large, diverse dataset means common backgrounds frequently appear with various objects or animals, such as cats, dogs, and elephants alongside blue skies, white clouds, houses, and trees. Therefore the model will learn background information unconsciously when doing classification tasks. Then fine-tuning on a smaller cat-dog dataset will yield better results because the background won’t interfere with the recognition of dog features. Without sufficient data, even a network with strong fitting capabilities cannot perform well and is prone to overfitting. We will consider this issue in Section 4.4.
Figure 5: OxfordIIITPet dataset’s accuracy on different classes.
Figure 6: Flower-102 dataset’s accuracy on different classes.
4.3 Model Depth Effect
Based on the conclusions from Section 4.2, this section explores the impact of network depth on prediction performance. According to equations (8) and (9), the depth of the model primarily affects the degrees of freedom of the model parameters within the UAT. Therefore, we designed different versions of the Para-Former model with 24 layers and depths of 2, 3, and 6, comparing their predictive performance against a ViT with a depth of 6. The results are shown in Table 3. The degrees of freedom for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are 48, 18, 72, 72, and 72, respectively (see Appendix B).
Why does Para-F-6-6, despite having the lowest degrees of freedom, only perform worse than Para-F-6-24? This is because the degrees of freedom encompass both the freedom of weights and biases. The freedom of weights can be understood as linear transformations derived from inputs, whereas the freedom of biases results from the approximation by multiple layers of UAT, which allows for stronger dynamic fitting capabilities. The deeper the layers, the stronger the approximation ability. Thus, Para-F-6-24 performs the best, followed by Para-F-6-6, as the number of UAT layer of their bias terms is largest (see Appendix B).
In summary, deeper models have stronger approximation capabilities, but there is a limit to how deep a model can go, necessitating a balance between network depth and layer count.
4.4 Dataset Effect
Cifar-10 Cifar-100 STL-10 OxfordIIITPet Flower-102
Para-F-12-12 98.62 91.36 99.47 91.87 96.89
Table 4: This table shows the results of fine-tuning using Google’s official pre-trained parameters of ViT. The pre-trained parameters were obtained from training the ViT model on the ImageNet dataset.
Figure 7: The fine-tuning process from the perspective of UAT.
After increasing the depth of the Para-Former, the results have improved to some extent. However, as shown in Table 3, almost all models encounter a performance bottleneck, especially on the OxfordIIITPet and Flower-102 datasets. What is the fundamental reason for this poor performance? This is likely because these datasets are relatively small and exhibit high data diversity. In this section, we will solve this problem.
First, the experimental results are based on validated ViT models (Para-F-1-1 and Para-Former-6-6 are equal to ViT with 1 and 6 depth.), whose performance is similar to that of the Para-Former. Given that ViT models are widely validated and are among the commonly used model architectures, their effectiveness is beyond doubt. So why is the performance so poor? The fundamental reason lies in the data itself: the model’s performance is typically directly related to the amount of available data. Take the Flower-102 dataset as an example, which contains 102 classes of flowers but only 1,020 training images, with some reserved for validation. This is far from sufficient for training deep learning models. Why is the performance even worse on the OxfordIIITPet dataset, which has a higher average number of images per class? This is due to the diversity of the data (i.e., the similarity of features within the same class) and the variability of the background.
Figures 5 and 6 show some results from the OxfordIIITPet and Flower-102 datasets. Both figures indicate that for categories with diverse backgrounds and inherent feature diversity, the prediction results are poor, as shown in Figures 5.b and 6.b. In contrast, when the background is consistent but the object morphology is diverse, the results are relatively better, as shown in Figures 5.a and 6.a. The best results are achieved when both the background and object morphology are minimal, as shown in Figures 5.c and 6.c. Although the Flower-102 dataset has fewer images, its backgrounds are more consistent, and the differences between species are smaller. Conversely, animals in the OxfordIIITPet dataset can appear in various forms, whereas flowers in the Flower-102 dataset are usually photographed from relatively fixed angles, leading to better performance.
Does this mean that deep learning models are not powerful enough or lack generalization and reasoning capabilities? Quite the opposite. With such limited data, deep learning models achieved about 25% accuracy on OxfordIIITPet and about 40% on Flower-102, actually demonstrating the strength and generalization ability of deep learning models. A common misconception about deep learning is that even if people are unfamiliar with certain species, by careful examination, they can accurately distinguish between OxfordIIITPet and Flower-102 images, leading to the belief that deep learning lacks generalization or human-like reasoning capabilities. This view is flawed. Suppose the human brain operates like a UAT model; it can automatically ignore background information before seeing these images because such backgrounds are frequently encountered in daily life, and the UAT in the brain has already fitted this information. Therefore, people only need to focus on specific information and then use this information to train the brain to identify similarities and differences. For example, a 10-year-old child’s brain, observing the world at 10 frames per second (Read and Meyer 2000), would have processed over $1.5\times 10^{9}$ high-resolution images by age 10 (Lu et al. 2017), many containing common objects and backgrounds from daily life. Hence, it is unfair to criticize deep learning models for being unable to train on small datasets and for lacking universality and generalization ability.
Therefore, we should train on large datasets and then fine-tune on smaller ones. This is also the current method for deep learning training, demonstrating the powerful generalization capabilities of deep learning models, especially when there is some correlation between the datasets. Following this approach, we fine-tuned a pre-trained ViT model using only 10 epochs. The results, as shown in Table 4, demonstrate significant improvements in prediction performance.
Figure 7 illustrates the rationale for fine-tuning. The goal, whether in training or prediction, is to obtain the target results based on the input. This can be understood as function fitting, expressed as $\mathbf{y}=\mathbf{G}(\mathbf{x}|\mathbf{A},\mathbf{W},\mathbf{\Theta})$, where $\mathbf{A}={\mathbf{\alpha}_{1}...\mathbf{\alpha}_{N}}$, $\mathbf{W}={\mathbf{W}_{1}...\mathbf{W}_{N}}$, and $\mathbf{\Theta}={\mathbf{\theta}_{1}...\mathbf{\theta}_{N}}$. The purpose of training is to determine $\mathbf{A},\mathbf{W},\mathbf{\Theta}$, which can dynamically change based on the input.
This means that if the background is similar but the objects are different, the resulting parameters will differ. If the model can make correct predictions, it implies that the model can distinguish between the background and the animals. If it could not be distinguished, it might incorrectly use the background as a feature of the animals, leading to the same classification. Moreover, if the model can correctly classify the animals, it indicates that the model can identify the features of the animals.
When the backgrounds are different, because natural backgrounds are generally similar, extensive training can differentiate background information. Thus, if the model can make right predictions, it suggests that the model has comprehensively grasped the animal features, even with significant environmental variations. Training on ImageNet follows this process. A new dataset is just a new series of combinations of various similar backgrounds and animals which have been seen in ImageNet. If the model has already learned sufficient background and animal information on ImageNet, it can dynamically fit the results based on the input. Therefore, only minor adjustments specific to new data are required during fine-tuning.
5 Conclusion
In this paper, we propose a parallel network based on the UAT and validate its effectiveness on large datasets. We analyze the characteristics of both serial and parallel networks. Serial networks based on residual connections, often have higher degrees of parameter freedom, leading to stronger generalization capabilities. However, as the network depth increases, the inference time also becomes longer. On the other hand, parallel networks trade computational resources for time, meaning their runtime is not affected by the number of network layers. This allows for the design of multi-layer parallel networks that can replace serial networks to achieve fast inference. Additionally, we analyze the impact and reasons data have on deep learning. From the UAT perspective, training on large datasets essentially enhances the model’s ability to fit diverse data.
Appendix A The UAT Format of Transformer-based ViTs
In this section, we will demonstrate that the mathematical form of single-layer and two-layer Transformers is equivalent to the UAT. Firstly, the basic form of a Transformer is shown in Figure 2 of main text. We represent the MHA and FFN in the Transformer in matrix-vector form as Eq. (10) and Eq. (11), respectively (for specific details, refer to UAT2LLMs). To distinguish between the operations and variables in the Transformer and those in the matrix-vector form, we add a prime () symbol to the operations and variables in the matrix-vector form. It is important to emphasize that the parameters of the MHA are influenced by the input.
$MHA^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (10)
$FFN^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (11)
Next, the result of a single-layer Transformer $\mathbf{x}^{\prime}_{1}$ is:
$\mathbf{x}^{\prime}_{1}$ $=\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{W}^{\prime}_{1,3}\sigma[\mathbf{W}^{\prime}_{1,2}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0})+\mathbf{b}^{\prime}_{1,2}]+\mathbf{b}^{\prime}_{1,3}$ (12)
$=\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{W}^{\prime}_{1,3}\sigma(\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})+\mathbf{b}^{\prime}_{1,3}$
Let $\mathbf{W}^{\prime}_{1,2}=\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}$, so we obtain:
$\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (13)
where $\mathbf{W}^{\prime}_{1,1}$ and $\mathbf{W}^{\prime}_{1,2}$ are parameters influenced by the input, thus having a degree of freedom of 2.
$\mathbf{x}^{\prime}_{2}$ $=\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})+\mathbf{b}^{\prime}_{2,3}$ (14)
$=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,3})+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})$
=$=\{\mathbf{W}^{\prime}_{2,1}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,3}\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,2}\}$
=$=\{\mathbf{W}^{\prime}_{2,1}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,3}+\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}({\mathbf{W}}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,2}+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
=$=\{(\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}})+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}\}$
In Eq. (14), we provide the matrix-vector format of 2-layer Transformer. Let $\mathbf{W}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}$, $\mathbf{b}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}}$, $\mathbf{W}^{\prime}_{1,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}$, $\mathbf{W}^{\prime}_{2,2}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, and $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$. Thus, Eq. (14) can be written as:
$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (15)
$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
where $\mathbf{W}^{\prime}_{2,1}$,$\mathbf{b}^{\prime}_{2,1}$,$\mathbf{W}^{\prime}_{1,2}$,$\mathbf{W}^{\prime}_{1,3}$,$\mathbf{W}^{\prime}_{2,3}$, and $\mathbf{b}^{\prime}_{2,2}$ are influenced by the input, so the degree of freedom is 6.
Another issue is that in $\mathbf{x}^{\prime}_{1}$ and $\mathbf{x}^{\prime}_{2}$, there is a term $(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})$ and $(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})$ which doesn’t generally conform to the standard terms in UAT. We can consider this term as a special case and regard it as a layer in UAT. Firstly, when the number of network layers increases, it effectively increases $N$ in UAT. Given UAT’s strong approximation capabilities, adding multiple instances of this term can be viewed as a linear transformation, thus it doesn’t affect the overall approximation power of UAT. Furthermore, this term can be considered a special form in UAT. Taking ReLU as an example activation function, $(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})$ can be expressed as $\text{ReLU}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3})-\text{ReLU}(-(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{1,3}))$.
In summary, we have shown that the mathematical forms of one-layer and two-layer Transformers can be seen as 2 and three-layer UAT.
Appendix B The Degree of Freedom of Different Layers Transformer
According to UAT2LLMs, the UAT format of multi-layer Transformer can be written as:
$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (16)
where $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$. The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by a $j$-layer of UAT with $\mathbf{x}_{0}$ as input. All $\mathbf{W}^{\prime}_{i,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input. Because they are affected by the weight of MHA in each layer. Why does the MHA operation affect the weight parameters in a linear manner? According to Eq. (12) and Eq. (13), this is because $\mathbf{W}^{\prime}_{1,2}=\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}$, where $\mathbf{W}^{\prime}_{1,1}$ represents the MHA parameters, which are influenced by the input, while $\mathbf{W}^{\prime}_{1,2}$ is a defined parameter. As a result, the parameter $\mathbf{W}^{\prime}_{1,2}$ in Eq. (13) becomes input-dependent, and this influence occurs through a linear transformation. Similarly, in a multi-layer Transformer, multiple layers of weight parameters are derived through similar transformations.
Therefore, it is easy to deduce that the degrees of freedom for Transformers with 1, 2, 3, and 6 layers are 2, 6, 9, and 18, respectively. For Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24, the degrees of freedom are shown in Table 1. At first glance, it seems that the degrees of freedom for Para-F-1-24, Para-F-2-24, Para-F-3-24, and Para-F-6-24 all exceed that of Para-F-6-6. Why then does only Para-F-6-24 consistently outperform Para-F-6-6?
Para-F-1-24 Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
Degree of Freedom 2*24 18*1  6*12 9*8 18*4
Table 1: The degree of freedom of Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, Para-F-6-24. We use the the degree of freedom of serial Transformer with the depth 1, 6, 2, and 3 to multiply the number of serial Transformer.
The reason lies in the nature of the degrees of freedom. Taking $\mathbf{W}^{\prime}_{2,2}$ and $\mathbf{b}^{\prime}_{2,2}$ from Eq. (15) as examples, $\mathbf{W}^{\prime}_{2,2}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, where $\mathbf{W}^{\prime}_{1,1}$ is an input-dependent parameter from MHA, and $\mathbf{W}^{\prime}_{2,2}$ is a defined parameter, becoming fixed post-training. Thus, the degrees of freedom for $\mathbf{W}^{\prime}_{2,2}$ in Eq. (15) are imparted through a linear transformation with the MHA parameter $\mathbf{W}^{\prime}_{1,1}$. Additionally, $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$. Hence, $\mathbf{b}^{\prime}_{2,2}$ in Eq. (15) can be understood as being approximated by a two-layer UAT, showcasing stronger dynamic transformation capabilities based on inputs. Due to UAT’s powerful capacity, $\mathbf{b}^{\prime}_{2,2}$ can dynamically fit specific functions based on inputs.
In a single-layer Transformer, as seen in Eq. (12), $\mathbf{b}^{\prime}_{1,3}$ and $\mathbf{b}^{\prime}_{1,2}$ are defined parameters, so the number of bias terms approximated using UAT is zero. For a two-layer Transformer, as shown in Eq. (15), the number of bias terms approximated using UAT is two, specifically $\mathbf{b}^{\prime}_{2,1}$ and $\mathbf{b}^{\prime}_{2,2}$, with the UAT layers approximating these bias terms being 1 and 2, respectively. According to Eq.(16), in a Transformer with $i+1$ layers, $\mathbf{b}_{i+1,1}$ is approximated by a single-layer UAT, while $\mathbf{b}^{\prime}_{j,2}$ is approximated by a UAT of $j$ layers for $j>1$. It is easy to deduce that for a three-layer Transformer, the number of bias terms approximated using UAT is three, with the UAT layers being 1, 2, and 3. For a six-layer Transformer, the number of bias terms approximated using UAT is six, with the UAT layers being 1, 2, 3, 4, 5, and 6. The UAT layers of bias terms for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are shown in Table 2.
Para-F-1-24 Para-F-6-6 Para-F-2-24 Para-F-3-24 Para-F-6-24
UAT Layers (0)*24 (1+2+3+4+5+6)*1 (1+2)*12 (1+2+3)*8 (1+2+3+4+5+6)*4
Table 2: The UAT layer count corresponding to the bias terms in Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 is calculated by summing the UAT layers associated with each bias term within a serial Transformer. Since each serial Transformer has multiple bias terms, with each term corresponding to different UAT layer counts, we represent the UAT layers by summing the UAT layers for each bias term. For example, ”1+2+3” indicates that three bias terms are approximated using UAT, with 1, 2, and 3 layers respectively. Additionally, since multiple serial Transformers are used in parallel, we multiply the sum by the number of parallel Transformers.
In summary, more layers mean more degrees of freedom in bias terms, leading to stronger model fitting capabilities. To achieve performance comparable to that of deep serial networks while ensuring model fitting ability and fast inference, parallel networks need to strike a balance between network depth and layer count. Parallel networks are designed for extremely large models. For instance, if a serial network with 1 million layers can be replaced by a parallel network with a depth of 10 and 100 million layers, the speedup factor can reach up to 100,000 times. Sure, we could improve the design of the Para-former to achieve speedup computing with fewer layers.